# Exploring the Key Features of Repeating Fast Radio Bursts with Machine Learning

### v2

@@ 3.4 Results Verification
As shown in Figure 2, there is one repeating FRB (FRB20181030A, represented with a red dot) falling within the non-repeater cluster. We compare the observational information and the dynamic spectrum images of FRB20181030A with its subsequent repeat, FRB20181030B, neither of which exhibits any sub-bursts. In Table 4, we provide a comparison of basic observational parameters, including sky localization, DM values, and other fundamental observables. While these bursts do not initially appear to originate from two distinct sources, given the localization uncertainties, we cannot entirely rule out this possibility. By inspecting their dynamic spectrum images, both bursts intuitively appear to be repeating bursts with narrow-band emission. The narrow-band feature of FRB20181030B (see https://www.chime-frb.ca/catalog/FRB20181030B) is particularly prominent. However, we note that FRB20181030A (see https://www.chime-frb.ca/catalog/FRB20181030A) has a low signal-to-noise ratio and its frequency spectrum is noise-dominated. This may have affected the reliability of its derived parameters, resulting in FRB20181030A being incorrectly fitted as broad-band emission. Therefore, we suspect that the classification of the repeater FRB20181030A as a non-repeater may be due to data quality issues rather than its intrinsic properties.
@@
TNS Name Repeater Name RA (J2000) Dec (J2000) DM $S_{\nu}$ $F_{\nu}$ $\gamma$ $r$ SNR
@@
To verify the reliability of our repeater candidates’ identification, we validate our predictions using an updated repeater catalog (Chime/Frb Collaboration et al., 2023). There are $25$ repeating FRBs reported and $6$ of them were previously observed as non-repeaters in the first CHIME/FRB Catalog. These six newly confirmed repeaters are marked with green stars in Figure 8.
@@
There is one newly confirmed repeater, i.e. FRB20180910A, located in the non-repeater cluster in the t-SNE embedding space, which is also marked with the green star. By analyzing the dynamic spectrum of FRB20180910A, we discover that this repeater exhibits a broadband emission characteristic similar to non-repeaters, which is unique compared to the narrowband emission typically observed in repeaters. To investigate that individual burst, such as FRB20180910A, do not significantly bias our results, we repeat the t-SNE embedding and clustering analysis after excluding this burst. The analysis confirmed that the removal of FRB20180910A did not alter the general clustering structure and the main conclusions of this study. This consistency demonstrates the robustness of our approach to outlier removal.
@@ 3.5 Parameters Distribution
The ECDFs and AD test results for the reclassified Sgrp sample (comprising 94 repeaters, 163 repeater candidates, and 337 non-repeaters) are displayed in Figure 9, with the solid and dashed red lines representing repeaters and non-repeaters, respectively. The AD test results are indicated by $p_{\mathrm{AD1}}$, which represents the statistical significance of the difference between the two sample distributions, and $s_{\mathrm{AD1}}$, the normalized test statistic quantifying the extent of this difference. Furthermore, to better investigate the differences in parameters distributions between repeating and non-repeating FRB sources, we conduct an additional analysis after excluding all sub-bursts from the sample. Specifically, for bursts with multiple sub-bursts, we retain only the event with sub_num = 0. Additionally, we compare only the first-detected repeater events for each repeating source, since these events were triggered at the nominal threshold, thus minimizing the statistical complications arising from multiple events per source. The resulting filtered sample includes 18 repeaters, 151 repeater candidates, and 317 non-repeaters. The ECDF analysis and AD tests for the sub-burst-excluded sample are also presented in Figure 9, with the solid and dashed blue lines representing repeaters and non-repeaters, respectively, and $p_{\mathrm{AD2}}$ and $s_{\mathrm{AD2}}$ showing the test results.
@@
where $\rm DM_{MW},DM_{Halo},DM_{IGM},$ and $\rm DM_{Host}$ denote the contributions from the Milky Way interstellar medium, the Milky Way halo, the intergalactic medium (IGM), and FRB host galaxy. $\rm DM_{MW}$ is estimated based on the NE2001 model (Cordes & Lazio, 2002). Therefore, the extragalactic DM can be defined as:
@@ Acknowledgments
Name Sub_num RA (J2000) Dec (J2000) DM $S_{\nu}$ $F_{\nu}$ $\Delta t_{\mathrm{WS}}$ $\Delta t_{\mathrm{ST}}$ $\gamma$ $r$ $\nu_{\mathrm{Low}}$

### v3

@@ 3.4 Results Verification
As shown in Figure 2, there is one repeating FRB (FRB20181030A, represented with a red dot) falling within the non-repeater cluster. We compare the observational information and the dynamic spectrum images of FRB20181030A with its subsequent repeat, FRB20181030B, neither of which exhibits any sub-bursts. In Table 4, we provide a comparison of basic observational parameters, including sky localization, DM values, and other fundamental observables. While these bursts do not initially appear to originate from two distinct sources, given the localization uncertainties, we cannot entirely rule out this possibility. By inspecting their dynamic spectrum images, both bursts intuitively appear to be repeating bursts with narrowband emission. The narrowband feature of FRB20181030B (see https://www.chime-frb.ca/catalog/FRB20181030B) is particularly prominent. However, we note that FRB20181030A (see https://www.chime-frb.ca/catalog/FRB20181030A) has a low SNR and its frequency spectrum is noise-dominated. This may have affected the reliability of its derived parameters, resulting in FRB20181030A being incorrectly fitted as broadband emission. Therefore, we suspect that the classification of the repeater FRB20181030A as a non-repeater may be due to data quality issues rather than its intrinsic properties.
@@
TNS Name Repeater Name R.A. (J2000) Decl. (J2000) DM $S_{\nu}$ $F_{\nu}$ $\gamma$ $r$ SNR
@@
To verify the reliability of our repeater candidates’ identification, we validate our predictions using an updated repeater catalog (Chime/Frb Collaboration et al., 2023). There are $25$ repeating FRBs reported and six of them were previously observed as non-repeaters in the first CHIME/FRB Catalog. These six newly confirmed repeaters are marked with green stars in Figure 8.
@@
There is one newly confirmed repeater, i.e. FRB20180910A, located in the non-repeater cluster in the t-SNE embedding space, which is also marked with the green star. By analyzing the dynamic spectrum of FRB20180910A, we discover that this repeater exhibits a broadband emission characteristic similar to non-repeaters, which is unique compared to the narrowband emission typically observed in repeaters. To investigate that individual bursts, such as FRB20180910A, do not significantly bias our results, we repeat the t-SNE embedding and clustering analysis after excluding this burst. The analysis confirmed that the removal of FRB20180910A did not alter the general clustering structure and the main conclusions of this study. This consistency demonstrates the robustness of our approach to outlier removal.
@@ 3.5 Parameters Distribution
The ECDFs and AD test results for the reclassified Sgrp sample (comprising 94 repeaters, 163 repeater candidates, and 337 non-repeaters) are displayed in Figure 9, with the solid and dashed red lines representing repeaters and non-repeaters, respectively. The AD test results are indicated by $p_{\mathrm{AD1}}$, which represents the statistical significance of the difference between the two sample distributions, and $s_{\mathrm{AD1}}$, the normalized test statistic quantifying the extent of this difference. Furthermore, to better investigate the differences in parameter distributions between repeating and non-repeating FRB sources, we conduct an additional analysis after excluding all sub-bursts from the sample. Specifically, for bursts with multiple sub-bursts, we retain only the event with sub_num = 0. Additionally, we compare only the first-detected repeater events for each repeating source since these events were triggered at the nominal threshold, thus minimizing the statistical complications arising from multiple events per source. The resulting filtered sample includes 18 repeaters, 151 repeater candidates, and 317 non-repeaters. The ECDF analysis and AD tests for the sub-burst-excluded sample are also presented in Figure 9, with the solid and dashed blue lines representing repeaters and non-repeaters, respectively, and $p_{\mathrm{AD2}}$ and $s_{\mathrm{AD2}}$ showing the test results.
@@
where $\rm DM_{MW},DM_{Halo},DM_{IGM},$ and $\rm DM_{Host}$ denote the contributions from the Milky Way interstellar medium, the Milky Way halo, the intergalactic medium (IGM), and FRB host galaxy. $\rm DM_{MW}$ is estimated based on the NE2001 model (Cordes & Lazio, 2002). Therefore, the extragalactic DM can be defined as
@@ Acknowledgments
This work uses unsupervised machine learning methods to identify 163 repeater candidates in the first CHIME/FRB catalog, with their observational parameters listed in Table 6. These candidates exhibit characteristics highly similar to repeaters, suggesting that they are more likely to repeat. Thus, they can be considered key observational targets for future studies on repeaters. Table 6 presents their positional information, including Right Ascension (R.A.), declination (Decl.), and DM, along with seven selected feature parameters describing the intrinsic or environmental properties of FRBs, i.e., $S_{\nu}$, $F_{\nu}$, $\Delta t_{\mathrm{WS}}$, $\Delta t_{\mathrm{ST}}$, $\gamma$, $r$ and $\nu_{\mathrm{Low}}$.
@@
Name Sub_num R.A. (J2000) Decl. (J2000) DM $S_{\nu}$ $F_{\nu}$ $\Delta t_{\mathrm{WS}}$ $\Delta t_{\mathrm{ST}}$ $\gamma$ $r$ $\nu_{\mathrm{Low}}$